Sol Summary – April 24th

Crew 297 Sol Summary Report 24-04-2024

Sol: 10

Summary Title: The Martian Chronicles

Author’s name: Pawel Sawicki

Mission Status: Nominal

Sol Activity Summary:

This morning a portion of the crew woke up early to evade the impending hot weather and venture out on an EVA to Somerville Overlook. With the morning breeze assisting, the crew took the brisk walk from Opportunity to the overlook and admired the views. But of course, research was the main priority, and as such the crew went to work. Many samples and measurements were collected from the area, and on the way back Case was finally able to return to areas outside of the MDRS station.

Meanwhile, the later EVA crew, within the warming weather, was able to finally find the hidden NPS and return it safely to the Hab. The crew of two then set foot to Skyline Rim and investigated NPS digging sites along the way, testing the Martian surfaces for a qualitative assessment of ease of digging. This particular EVA broke a record for longest walking distance during Mission 297 – a Brobdingnagian 10 km!

Look Ahead Plan:

Tomorrow will be the last day involving multiple EVAs. The morning EVA will venture out to the Gateway of Lith and take Case out for its last stroll into Martian regolith. Case is expected to now be fully functional and should reliably take imagery of the area (with some expected intervention via crew manual driving). The second EVA will venture out to Somerville Overlook to scope areas for NPS digging. Back at the MDRS station, the crew will prepare the decommissioning process of the GreenHab for the season.

Anomalies in work:

Suit 11 (fan power substantially decreased mid-EVA): Battery was replaced but the air flow is still less than nominal. Inop’d for the current mission.
Morning (data from 0900): calm winds (3 mph), chilly (59F), and sunny with clear skies
Noon: moderate winds (12 mph), warm (76F), and sunny with clear skies
Afternoon (1800): moderate winds (14 mph), hot (81F), and sunny with clear skies
Crew Physical Status: Overall crew health is nominal.

EVA: There were two EVAs today. The first EVA, involving Sarah and Pawel, successfully reached the Somerville Overlook and took gamma-ray spectrometer measurements along with samples. On the way back, near Compass Rock, the crew performed the human-interaction study with Case, though had to cut short due to time constraints. The second EVA was able to (finally) successfully locate the NPS and bring it back to the Hab safely. The crew also ventured to Skyline Rim and located many potential NPS digging locations on the way.

Reports to be filed:

Sol Summary
Journalist Report
GreenHab Report
Operations Report
EVA report (2 EVAs)
EVA request (2 EVAs)
EVALink Feedback Report
Support Requested: None requested.

Copyright © The Mars Society. All rights reserved. | Main Site